OpenStudy (anonymous):

How do you make a sine function become a cosine function? Like I have f(x)=14.25 sin(0.08πx)+15.75, but I need to make it into a cosine function and I'm not sure how I should do that...?

OpenStudy (mathstudent55):

\(\cos \theta = \sin (90 - \theta) \)

OpenStudy (anonymous):

Oh, why not sin(x) = cos(π/2 - x) ?

OpenStudy (mathstudent55):

Since you have sine in your problem, and you need cosine, take the argument of the sine and subtract it from 90 or \(\pi /2\), and you will have transformed it into the cosine. The way you wrote it seems like you want to transform a cosine into a sine.

OpenStudy (anonymous):

Ohh alright makes sense, thank you!

OpenStudy (mathstudent55):

wlcm

OpenStudy (anonymous):

An easy way to remember the transforms is to look at the graph of sine and cosine. They are identical with a phase shift of 90 degrees (or pi/2 radians)